# Learning Top-$k$ Subtask Planning Tree based on Discriminative Representation Pre-training for Decision Making

### v1

@@ 3.3. Top-$k$ Subtask Planning Tree Generation
where $W_{\phi},W_{\varphi}$ are learnable parameters and $\beta_{q,i}$ is the attention score. Here, we can model the correlation between query vector $v^{q}$ and an arbitrary number of subtasks, which allows us to easily increase the number of subtasks in our tree generation process. Then the categorical distribution can be modeled using the weights from which the variable $C$ can be sampled:
@@
Then, we use the sampled top-$k$ subtask ids as the $k$ nodes at a certain layer in the subtask tree (Please note that we set $k=1$ while generating the root node). Moreover, given the hyperparameter $N$, we can obtain the tree of depth $N$ by repeating the above process based on the planning transition (described in the next paragraph), and the original correlation weight ${\alpha_{q,i}}$ is set as connections (edges) between layer and layer. For a clear understanding, we summarize the overall construction procedure shown in figure 4.
Planning for future: $m$-step predictor
While expanding the subtask tree, we should make planning for the $m$-step prediction of the future. We design a $m$-step predictor denoted as $p_{m}(\hat{s}_{t+m}|\hat{s}_{t},\hat{a}_{t})$ that is used to derive the next $m^{th}$ step’s state. The action $\hat{a}_{t+m}$ is output by the policy network $\pi$ (defined in Equation (16)) using the current subtask representation. The $N$-times rollout starting with $s_{t-1}$ can be written as:
$\left\{{({s_{t-1}},{a_{t-1}}),({{\hat{s}}_{t+m-1}},{{\hat{a}}_{t+m-1}}),...,({{\hat{s}}_{t+N\cdot m-1}},{{\hat{a}}_{t+N\cdot m-1}})}\right\}.$
Then we can use this imagined rollout to expand the subtask tree to a greater depth. That is, as each layer of depth increases, the planning will perform $m$-step imagination forward.
Figure 4. The basic top-$k$ subtask tree construction process.
Optimization for tree generation
The optimization includes two parts: an $m$-step predictor $p_{m}$ and an attention module. The former $p_{m}$ can be trained by minimizing the mean squared error function $\mathcal{L}_{m}$ defined as follows. Sampled some state-action pairs from replay buffer $\mathcal{D}$, we have:
$\mathcal{L}_{m}={\mathbb{E}_{\scriptstyle({s_{t}},{a_{t}},{s_{t+m}})\sim\mathcal{D}\hfill\atop\scriptstyle{{\hat{s}}_{t+m}}={p_{m}}({s_{t}},{a_{t}})\hfill}}\left[{||{{\hat{s}}_{t+m}},{s_{t+m}}||_{2}^{2}}\right].$ (10)
The attention module parameterized by $\omega=\{\phi,\varphi\}$ can be optimized in a Monte-Carlo policy gradient manner Williams (1992). First, we define the intra-subtask cumulative reward $R_{t}=\sum\nolimits_{j=0}^{\tilde{T}-1}{{\gamma^{j}}{r_{t+j}}}$ during the interval $\tilde{T}$ of the subtask execution process as follows. By decomposing the cumulative rewards into a Bellman Expectation equation in a recursive manner, we acquire:
@@
where ${s_{t}}$ and ${\mathcal{T}_{t}}$ is the current state and the executed subtask, $\gamma$ is the discounted factor, and ${\mathcal{T}_{t+\tilde{T}}}$ is the next subtask.
@@
3.4. Tree-auxiliary Policy
Given a generated top-$k$ subtask tree, we can obtain a tree-auxiliary policy as follows. Generally, agents should have the foresight to judge the current situation, give a rational plan, and then use the plan to guide policy learning. Therefore, the first step is to generate a plan based on the top-$k$ subtask tree.
Subtask execution plan generation
Here, we design a discounted upper confidence bound (d-UCB) method for the path finding. For each subtask $\mathcal{T}^{i}$, we record its selection count $c_{\mathcal{T}^{i}}$ and the average cumulative rewards $r_{\mathcal{T}^{i}}$ while the subtask terminates. The UCB value for the $\mathcal{T}^{i}$ in the planning tree is computed as:
@@ 3.4. Tree-auxiliary Policy
Then, we give a definition of the discounted path length via the evaluation of d-UCB. We adopt the greedy policy to pick the path (subtask execution plan) with the maximum discounted length. The detailed intuition description is shown in Appendix A.2.
@@
Sample a path from root to leaf in the tree, the length of the discounted path can be defined as follows.
${{\hat{d}}}(root\to leaf)=\sum\nolimits_{l}{{\kappa^{l}}{e^{l,l+1}}\cdot v_{\mathcal{T}^{i}}},$
where $e^{l,l+1}$ is a sampled edge between $l^{th}$ layer and $(l+1)^{th}$ layer, the value of the edge is the probability of subtask transition, $\kappa=0.8$ is the discounted factor, and $v_{\mathcal{T}^{i}}$ is the UCB value of the sampled node in the $(l+1)^{th}$ layer.
Subtask termination condition
Subtask execution should be terminated when the current situation changes, and we give a measurement about when to terminate the current subtask. we first calculate the cosine similarity of the current query vector $\mathcal{E}^{q}(s_{t},a_{t})$ and the corresponding subtask embedding $\mathcal{E}^{\mathcal{T}_{id}}(s_{t},a_{t})$ and the similarity function is defined as follows.
@@
For every timestep, the agent will compare the similarity score of the current timestep and the last timestep. If the similarity score decreases more than a threshold $\Delta=0.5$, the subtask will terminate and turn to the next one until the path ends.
Subtask-guided decision making
In our implementation, the policy will consider the current state and the subtask embedding and give a comprehensive decision, formulated as follows.
@@
As for the policy optimization, we adopt the PPO-style Schulman et al. (2017) training regime and the objective function is defined as follows.
${\mathcal{L}_{\pi}}=\left[{\min\left({{i_{t}}\cdot{A_{t}},clip({i_{t}},1-\varepsilon,1+\varepsilon){A_{t}}}\right)}\right],$ (16)
where $i_{t}=\frac{{\pi({a_{t}}|{s_{t}},{{{{\mathcal{E}}}}^{{{{{\mathcal{T}}}}_{id}}}}(s_{t-1},a_{t-1}))}}{{{\pi_{old}}({a_{t}}|{s_{t}},{{{{\mathcal{E}}}}^{{{{{\mathcal{T}}}}_{id}}}}(s_{t-1},a_{t-1}))}}$ is the importance ratio, and $A_{t}$ is is an estimator of the advantage function at timestep $t$.
@@
All in all, the overall optimization flow including subtask representation pre-training, top-k subtask generation and tree-auxiliary policy is summarized in Algorithm 1.
Ensure subtask representation policy $\rho=\{\{\mathcal{E}^{i}\}_{i=1}^{n},f_{pre}\}$, tree building policy $\mathcal{E}^{q}$, and tree-axuliary policy $\pi$;
// Subtask Representation Pre-training
Collect diverse subtask datasets $\{\mathcal{T}^{1},...,\mathcal{T}^{k}\}$;
for each iteration do
       Sample a subtask $\mathcal{T}^{i}$ and transitions $\{\tau^{i}_{t}\}_{t=1}^{b}$, $\{\tau^{i,+}_{t}\}_{t=1}^{b}$;
       $\{z^{i}_{t}\}_{t=1}^{l}=\{\mathcal{E}^{i}(s_{t}^{i},a_{t}^{i})\}_{t=1}^{l}$ ;
       $\{z^{i,+}_{t}\}_{t=1}^{l}=\{\mathcal{E}^{i}(s_{t}^{i,+},a_{t}^{i,+})\}_{t=1}^{l}$ ;
       Sample other subtasks $\mathcal{T}\backslash\mathcal{T}^{i}$ and transitions $\{\tau^{i,-}_{t}\}_{t=1}^{b}$;
       $\{z^{i,-}_{t}\}_{t=1}^{l}=\{\mathcal{E}^{-}(s_{t}^{i,-},a_{t}^{i,-})\}_{t=1}^{l}$ ;
       Predict reward $r$ and next state $s^{\prime}$;
       Update $\rho$ to minimize Eq. (4);
end for
// Tree Generation and Tree-auxiliary Policy
for each episode do
       Initial state-action pair $s\_a_{cur}=(s_{0},a_{0})$;
       BUILD_TREE($s\_a_{cur},\mathcal{E}^{q},\{\mathcal{E}^{i}\}_{i=1}^{n}$) ; // Fig. 4
       Use d-UCB in Sec. 3.4 to find a plan $\zeta$;
       for each timestep $t$ do
             if  $\zeta$ is None  then
                   Regenerate the tree and the plan;
             end if
            while $\zeta$ is not None do
                   $\mathcal{T}_{id}=\zeta[0]$;
                   // execute current subtask Select action $a_{t}=\pi(\cdot|s_{t},\mathcal{E}^{\mathcal{T}_{id}}(s\_a_{cur}))$;
                   Receive reward $r$ and next state $s^{\prime}$ ;
                   if  $Sim({{{{\mathcal{E}}}}^{q}}({s_{t}},{a_{t}}),{{{{\mathcal{E}}}}^{{{{{\mathcal{T}}}}_{id}}}}({s_{t}},{a_{t}}))<0.5$  then
                         remove $\mathcal{T}_{id}$ from $\zeta$; // execute next subtask
                   end if
             end while
       end for
      Update $\pi$ with Eq. (16);
       Update $\mathcal{E}^{q}$ with Eq. (10) and Eq. (12);
end for
Algorithm 1 Overall Procedure
4. Experiments
@@
In this section, we seek to empirically evaluate our method to verify the effectiveness of our solution. We design the experiments to 1) verify the effectiveness of the multiple-encoder and individual-predictor learning regime; 2) test the performance of our top-$k$ subtask tree on the most complicated scenarios in the 2D navigation task BabyAI Chevalier-Boisvert et al. (2019); 3) conduct some ablation studies in our settings.
@@
Figure 5. The schematics of four basic subtasks. The dashed lines denote the path to reach the goal. The agent represented by the red triangle is partially observable to the grid and the light-grey shaded area represents its field of view. The missions are described as : (a). Open the red door; (b). Go to the red box; (c). Put the grey box next to the red key. (d). Pick up the green ball.
Figure 6. The illustration for one of the BossLevel scenario. The mission is: ”pick up a red box and go to the purple door”.
4.1. Experimental Setting
Basic Subtask
@@ 4.1. Experimental Setting
BabyAI contains four basic subtasks that may appear in more complicated scenarios, including Open, Goto, PutNext, and Pickup, as shown in figure 5. (a) Open : Open the door with a certain color and a certain position, e.g. ”open the green door on your left”. The color set and position set are defined as : $COLOR=\{red,\ green,$ $blue,\ purple,$ $yellow,\ grey\}$ and $POS=\{on\ your\ left,\ on\ your\ right,$ $\ in\ front\ of\ you,\ behind$ $\ you\}$. (b) Goto : Go to an object that the subtask instructs, e.g. ”go to red ball” means going to any of the red balls in the maze. The object set is defined as: $OBJ=\{door,\ box,\ ball,\ key\}$. (c) PutNext : Put objects next to other objects, e.g. ”put the blue key next to the green ball”. (d) PickupLoc : Pick up some objects, e.g.“pick up a box”.
Complicated Task
@@
The BabyAI platform includes some tasks that require three or more subtasks compositions to solve, which are called complicated tasks. What we want to do is to solve these tasks through our tree-based policy to reach higher performance and faster convergence. Take the ”BossLevel” scenario for example, the agent should follow one of the instructions like ”pick up the grey box behind you, then go to the grey key and open a door”, which contains three subtasks and is shown as figure 6. More detailed descriptions are shown in Appendix A.3.
Baselines
In this paper, we select the state-of-the-art (SOTA) algorithm PPO Schulman et al. (2017) and SAC Haarnoja et al. (2018), the tree-based planning algorithm SGT Jurgenson et al. (2020), and the context-based RL method CORRO Yuan and Lu (2022) as baselines to verify the effectiveness of our proposed method. PPO is a popular deep reinforcement learning algorithm for solving discrete and continuous control problems, which is an on-policy algorithm. SAC is an off-policy actor-critic deep RL algorithm based on the maximum entropy reinforcement learning framework. SGT adopts the divide-and-conquer style that recursively predicts intermediate sub-goals between the start state and goal. It can be interpreted as a binary tree that is most similar to our settings. CORRO is based on the framework of context-based meta-RL, which adopts a transition encoder to extract the latent context as a policy condition.
Parameter Settings
@@
Additionally, we provide the description of all the neural networks used in our framework and more detailed experimental hyperparameter settings in Appendix A.4 and Appendix A.5, respectively.
4.2. Results on Subtask Representaion
Here report the experimental visualization results on the settings described in Section 3.2 to verify the effectiveness of our distinguishable subtask representations, shown as Figure 7.
@@ 4.2. Results on Subtask Representaion
Figure 7. The t-SNE visualization of subtask embeddings. (a-c): The results are generated by our multiple encoders and one predictor learning regime on 2, 3, and 4 subtasks, respectively. (d-f): Substitute our multiple encoder modules with one shared encoder and regenerate the visualization.
Firstly, we train our multiple encoders and individual predictor with collected diverse subtask datasets on 2, 3, and 4 subtasks, respectively. Then we substitute the multiple encoder modules with one shared encoder and train it. Figure 7 visualizes the t-SNE Van der Maaten and Hinton (2008) 2D projection of the latent subtask embeddings for these two models. The results show that our regime is capable of clearly illustrating separate clusters of embeddings for different tasks no matter how many subtasks or encoders we use. However, only one shared encoder can not learn the boundaries and distinguishment clearly and we speculate that mixed representation training for multiple different subtasks would interfere with the stabilization of the representation module.
Moreover, due to the powerful capacity of contrastive learning and model-based dynamic prediction, our model can exhibit strong representational discrimination and contain enough dynamics knowledge, which will be verified empirically as follows.

### v2

@@ 4.3 Top-K Subtask Planning Tree Generation
where $W_{\phi},W_{\varphi}$ are learnable parameters and $\beta_{q,i}$ is the attention score. Here, we can model the correlation between query vector $v^{q}$ and an arbitrary number of subtasks, which allows us to easily increase the number of subtasks in our tree generation process. Then the categorical distribution can be modelled using the weights from which the variable $C$ can be sampled:
@@
Then, we use the sampled top-$K$ subtask ids as the $K$ nodes at a certain layer in the subtask tree (please note that we set $k=1$ while generating the root node). Moreover, given the hyperparameter $N$, we can obtain the tree of depth $N$ by repeating the above process based on the planning transition (described in the next paragraph), and the original correlation weight ${\alpha_{q,i}}$ is set as connections (edges) between layers. For a clear understanding, we summarize the pseudo-code shown in Fig. 3b.
Planning for future: $m$-step predictor. While expanding the subtask tree, we should make planning for the $m$-step prediction of the future. We design an $m$-step predictor denoted as $p_{m}(\hat{s}_{t+m}|\hat{s}_{t},\hat{a}_{t})$ that is used to derive the next $m^{th}$ step’s state. The action $\hat{a}_{t+m}$ is output by the policy network $\pi$ (defined in Equation (16)) using the current subtask representation. The $N$-times rollout starting with $s_{t-1}$ is as follows:
$\left\{{({s_{t-1}},{a_{t-1}}),({{\hat{s}}_{t+m-1}},{{\hat{a}}_{t+m-1}}),...,({{\hat{s}}_{t+N\cdot m-1}},{{\hat{a}}_{t+N\cdot m-1}})}\right\}.$ (9)
Then we can use this imagined rollout to expand the tree to a greater depth. That is, as each layer of depth increases, the planning will perform $m$-step imagination.
Optimization for tree generation. The optimization includes two parts: an $m$-step predictor $p_{m}$ and an attention module. The former $p_{m}$ can be trained by minimizing the mean squared error function $\mathcal{L}_{m}$ defined as follows. By sampling some state-action pairs from replay buffer $\mathcal{D}$, we have:
$\mathcal{L}_{m}={\mathbb{E}_{\scriptstyle({s_{t}},{a_{t}},{s_{t+m}})\sim\mathcal{D}\hfill\atop\scriptstyle{{\hat{s}}_{t+m}}={p_{m}}({s_{t}},{a_{t}})\hfill}}\left[{\|{{\hat{s}}_{t+m}},{s_{t+m}}\|_{2}^{2}}\right].$ (10)
The attention module formulated by Equations (6), (7), and (8), and parameterized by $\omega=\{\phi,\varphi\}$ can be optimized in a Monte-Carlo policy gradient manner williams1992simple . First, we define the intra-subtask cumulative reward $R_{t}=\sum\nolimits_{j=0}^{\tilde{T}-1}{{\gamma^{j}}{r_{t+j}}}$ during the interval $\tilde{T}$ of the subtask execution process as follows. By decomposing the cumulative rewards into a Bellman Expectation equation in a recursive manner, we acquire:
@@
where ${s_{t}}$ and ${\mathcal{T}_{t}}$ are the current state and the executed subtask, respectively. $\gamma$ is the discounted factor, and ${\mathcal{T}_{t+\tilde{T}}}$ is the next subtask.
@@
In summary, our top-$K$ subtask planning tree module revolutionizes decision-making by dynamically creating subtask trees to assist in policy training. Inspired by human-like foresight, we adjust the subtask tree during episodes, rather than relying on predefined structures, effectively handling long horizons and sparse feedback in complex, dynamic tasks.
4.4 Tree-auxiliary Policy
To obtain a tree-auxiliary policy, we start by generating a plan from the top-$K$ subtask tree. This plan serves as a guideline for decision-making, enabling agents to make rational actions based on a foresight judgment of the situation.
Subtask execution plan generation. We propose a discounted upper confidence bound (d-UCB) method for pathfinding, where we record the selection count $c_{\mathcal{T}^{i}}$ and the average cumulative rewards $r_{\mathcal{T}^{i}}$ for each subtask $\mathcal{T}^{i}$ while terminating. The UCB value for the $\mathcal{T}^{i}$ in the planning tree is computed as:
@@ 4.4 Tree-auxiliary Policy
Then, we define the discounted path length using d-UCB and then use the greedy policy to select the path with the maximum discounted length. More details behind the d-UCB design can be found in Appendix A.
@@
Sample a path from the root to leaf in the tree, the length of the discounted path can be defined as follows.
${{\hat{d}}}(root\to leaf)=\sum\nolimits_{l}{{\kappa^{l}}{e^{l,l+1}}\cdot v_{\mathcal{T}^{i}}},$ (14)
where $e^{l,l+1}$ is a sampled edge between the $l^{th}$ layer and $(l+1)^{th}$ layer, the value of the edge is the probability of subtask transition, $\kappa=0.8$ is the discounted factor, and $v_{\mathcal{T}^{i}}$ is the UCB value of the sampled node in the $(l+1)^{th}$ layer.
Subtask termination condition. Subtask execution should be terminated when the current situation changes, and we provide a measurement about when to terminate the current subtask. We first calculate the cosine similarity of the current query vector $\mathcal{E}^{q}(s_{t},a_{t})$ and the corresponding subtask embedding $\mathcal{E}^{\mathcal{T}_{id}}(s_{t},a_{t})$ and the similarity function is defined as follows.
@@
For every timestep, the agent will compare the similarity score of the current timestep and the last timestep. Avoiding the drastic changes in subtasks, we choose $\Delta=0.5$ as the division. A similarity score of less than $0.5$ indicates a significant change in the environment, and the current subtask should terminate.
Algorithm 1 Overall Algorithm Procedure
1:subtask representation policy $\rho=\{\{\mathcal{E}^{i}\}_{i=1}^{n},f_{pre}\}$, tree building policy $\mathcal{E}^{q}$, and tree-axuliary policy $\pi$;
2:// Subtask Representation Pretraining
3:Collect diverse subtask datasets $\{\mathcal{T}^{1},...,\mathcal{T}^{n}\}$;
4:for each iteration do
5:     Sample a subtask $\mathcal{T}^{i}$ and transitions $\{\tau^{i}_{t}\}_{t=1}^{b}$, $\{\tau^{i,+}_{t}\}_{t=1}^{b}$;
6:     $\{z^{i}_{t}\}_{t=1}^{l}=\{\mathcal{E}^{i}(s_{t}^{i},a_{t}^{i})\}_{t=1}^{l}$ ;
7:     $\{z^{i,+}_{t}\}_{t=1}^{l}=\{\mathcal{E}^{i}(s_{t}^{i,+},a_{t}^{i,+})\}_{t=1}^{l}$ ;
8:     Sample other subtasks $\mathcal{T}\backslash\mathcal{T}^{i}$ and transitions $\{\tau^{i,-}_{t}\}_{t=1}^{b}$;
9:     $\{z^{i,-}_{t}\}_{t=1}^{l}=\{\mathcal{E}^{-}(s_{t}^{i,-},a_{t}^{i,-})\}_{t=1}^{l}$ ;
10:     Predict reward $r$ and next state $s^{\prime}$;
11:     Update $\rho$ to minimize Eq. (4);
12:end for
13:// Tree Generation and Tree-auxiliary Policy
14:for each episode do
15:     Initial state-action pair $s\_a_{cur}=(s_{0},a_{0})$;
16:     BUILD_TREE($s\_a_{cur},\mathcal{E}^{q},\{\mathcal{E}^{i}\}_{i=1}^{n}$) ; // Fig. 3b
17:     Use d-UCB in Sec. 4.4 to find a plan $\zeta$;
18:     for each timestep $t$ do
19:         if  $\zeta$ is None then
20:              Regenerate the tree and the plan;
21:         end if
22:         while $\zeta$ is not None do
23:              $\mathcal{T}_{id}=\zeta[0]$;
24:              // execute current subtask
25:              Select action $a_{t}=\pi(\cdot|s_{t},\mathcal{E}^{\mathcal{T}_{id}}(s\_a_{cur}))$;
26:              Receive reward $r$ and next state $s^{\prime}$ ;
27:              if  $Sim({{{{\mathcal{E}}}}^{q}}({s_{t}},{a_{t}}),{{{{\mathcal{E}}}}^{{{{{\mathcal{T}}}}_{id}}}}({s_{t}},{a_{t}}))<0.5$  then
28:                  remove $\mathcal{T}_{id}$ from $\zeta$; // execute next subtask
29:              end if
30:         end while
31:     end for
32:     Update $\pi$ with Eq. (16);
33:     Update $\mathcal{E}^{q}$ with Eq. (10) and Eq. (12);
34:end for
Subtask-guided decision making. In our implementation, the policy will consider the current state and the subtask embedding and give a comprehensive decision, formulated as follows.
@@
As for the policy optimization, we adopt the PPO-style schulman2017ppo training regime. We summarized the objective function as follows.
${\mathcal{L}_{\pi}}=\left[{\min\left({{i_{t}}\cdot{A_{t}},clip({i_{t}},1-\epsilon,1+\epsilon){A_{t}}}\right)}\right],$ (17)
where $i_{t}=\frac{{\pi({a_{t}}|{s_{t}},{{{{\mathcal{E}}}}^{{{{{\mathcal{T}}}}_{id}}}}(s_{t-1},a_{t-1}))}}{{{\pi_{old}}({a_{t}}|{s_{t}},{{{{\mathcal{E}}}}^{{{{{\mathcal{T}}}}_{id}}}}(s_{t-1},a_{t-1}))}}$ represents the importance ratio between the current policy $\pi$ and previous policy $\pi_{old}$, which corresponds to the policy employed in the last training step. $A_{t}$ is an estimator of the advantage function at timestep $t$. The parameter $\epsilon=0.2$ serves as the clipping threshold in the PPO algorithm. The clip($\cdot$) function is utilized to constrain its value of $i_{t}$ within the interval $[1-\epsilon,1+\epsilon]$. This function serves a critical role in mitigating excessively aggressive updates to the policy, thereby contributing to the overall stability and efficiency of the algorithm.
@@
The overall optimization flow including subtask representation pretraining, top-$K$ subtask generation, and tree-auxiliary policy is summarized in Algorithm 1.
In summary, the tree-auxiliary policy leverages the top-$K$ subtask planning tree generated through the dynamic decision-making process. This plan serves as a valuable guideline for agents, allowing them to make informed decisions based on forward-looking assessments of the current scenario. Incorporating subtask representations into the policy conditions broadens the agent’s perspective, enabling effective policy adjustments. This tree-auxiliary policy component complements our framework, contributing to its adaptability and robust decision-making capabilities.
4.5 Overall Analysis of Proposed Modules
In this subsection, we will elaborate on how our proposed Top-$K$ Subtask Planning Tree can tackle long-horizon missions and how the combination of subtask representations and the planning tree contributes to solving this particular challenge.
Question 1: Why do top-$K$ subtask planning tree resolve long-horizon missions?
In long-horizon tasks where agents must plan and execute a series of actions, traditional reinforcement learning methods face a hard challenge as they typically struggle with sparse feedback and signals spanning extensive time steps. Our top-$K$ subtask planning tree addresses this by fragmenting larger tasks into smaller, more manageable subtasks. Each subtask operates within a shorter timespan, receiving more frequent and explicit feedback, significantly simplifying the learning and decision-making process.
Question 2: What is the role of subtask representations?
Subtask representations enable our model to understand and handle tasks in a structured manner. By encoding each subtask, the model can more easily discern relationships and dependencies among tasks, assisting in planning and executing actions more effectively. This approach also offers a pathway for more efficient reuse of previously learned knowledge and skills, accelerating the learning process.
Question 3: What is the contribution of the top-$K$ subtask planning tree?
The planning tree further augments the capabilities of our approach. By utilizing a planning tree, our model can make both long-term and short-term decisions within a unified framework. At each node, the model dynamically generates and adjusts the subtask tree based on the current state and contextual information, achieving flexible and efficient decision-making. The planning tree also affords the model a means to understand and explain its decisions and actions, enhancing its interpretability and transparency.
Question 4: Why do we need the combination of these modules?
In summary, our top-$K$ subtask planning tree, by integrating subtask representations and a planning tree, effectively addresses the challenges posed by long-horizon tasks. Subtask representations facilitate the model in handling complex tasks with ease, while the planning tree provides a unified framework for efficient and flexible decision-making. The combination of these elements not only empowers our model to successfully accomplish long-horizon tasks but also maintains its explainability and transparency, enhancing its reliability and practicality in real-world applications.
5 Experiments
@@
In this section, we seek to empirically evaluate our method to verify the effectiveness of our solution. We design the experiments to 1) verify the effectiveness of the multiple-encoder and individual-predictor learning regime; 2) test the performance of our top-$k$ subtask tree on the most complicated scenarios in the 2D navigation task BabyAI babyai_iclr19 ; 3) conduct some sensitivity analyses in our settings; 4) give some intuitive statistics about our model.
5.1 Experimental Setting
@@
Figure 4: The schematics of four basic subtasks. The dashed lines denote the path to reach the goal. The agent represented by the red triangle is partially observable to the grid and the light-grey shaded area represents its field of view. The missions are described as : (a). Open the red door; (b). Go to the red box; (c). Put the grey box next to the red key. (d). Pick up the green ball.
@@
Basic Subtask. BabyAI contains four basic subtasks that appear in more complicated scenarios, including: (a) Open: Open the door with a certain color and a certain position, e.g. ”open the green door on your left”. The color set and position set are defined as : $COLOR=\{red,\ green,$ $blue,\ purple,$ $yellow,\ grey\}$ and $POS=\{on\ your\ left,\ on\ your\ right,$ $\ in\ front\ of\ you,\ behind$ $\ you\}$. (b) Goto: Go to an object that the subtask instructs, e.g. ”go to red ball” means going to any of the red balls in the maze. The object set is defined as: $OBJ=\{door,\ box,\ ball,\ key\}$. (c) PutNext: Put objects next to other objects, e.g. ”put the blue key next to the green ball”. (d) PickupLoc: Pick up some objects, e.g.“pick up a box”. The detailed subtask illustrations are shown in Fig. 4.
Figure 5: Illustration of one of the BossLevel scenarios. The mission is: “Pick up a red box and go to the purple door”.
@@
Complex Task. The BabyAI platform includes some tasks that require three or more subtask compositions to solve, which are called complicated tasks. What we want to do is to solve these tasks through our tree-based policy to reach higher performance and faster convergence. Take the ”BossLevel” scenario as an example, the agent should follow one of the instructions ”pick up the grey box behind you, then go to the grey key and open a door”, which contains three subtasks, as shown in Fig. 5. More detailed descriptions are shown in Appendix B.
Baselines. In this paper, we select the state-of-the-art (SOTA) algorithms PPO schulman2017ppo and SAC haarnoja2018sac , the tree-based planning algorithm SGT jurgenson2020sub , the context-based RL method CORRO yuan2022corro and DGRL islam2022discrete , and LM-based planner LEAP chen2023planning as baselines to verify the effectiveness of our proposed method.
  • PPO is a popular deep reinforcement learning algorithm for solving discrete and continuous control problems, which is an on-policy algorithm.
  • SAC is an off-policy actor-critic deep RL algorithm based on the maximum entropy reinforcement learning framework.
  • SGT adopts the divide-and-conquer style that recursively predicts intermediate subgoals between the start state and goal. It can be interpreted as a binary tree that is most similar to our settings.
  • CORRO is based on the framework of context-based meta-RL, which adopts a transition encoder to extract the latent context as a policy condition.
  • DGRL learns a latent representation for both observations and goals and converts it into discrete latent based on a vector quantized variational autoencoder quantization bottleneck for downstream RL tasks in the context of goal-conditioned hierarchical reinforcement learning.
  • LEAP utilizes a masked language model to iteratively plan for the decision-making process, leading to improved RL performance across different tasks.
To ensure the impartiality of our experiments, we pretrain the transition encoder for CORRO using the same subtask transitions that we collected for our own setting. Additionally, the dataset used for LEAP is identical to ours. We have included a table outlining key hyperparameter settings for the baseline models in Table 2, and for more detailed configurations, we refer readers to the original papers.
Algorithm Description Value
clipping parameter $0.1$
VF coeff. $c_{1}$ $1$
PPO Entropy coeff. $c_{2}$ $0.01$
GAE parameter $\lambda$ $0.95$
learning rate $10^{-4}$
SAC target smoothing coefficient $\tau$ $0.005$
learning rate $3*10^{-4}$
SGT depth $D$ $3$
learning rate $0.005$
negative pairs number $64$
CORRO task embedding size $5$
learning rate $3*10^{-4}$
codebook size $L$ $10$
DGRL the number of factors per vector $G$ $16$
task embedding size $5$
learning rate $3*10^{-4}$
embedding size $128$
Instruction Encoder channels $128$
LEAP State Encoder channels [256,256,128]][256,256,128]][ 256 , 256 , 128 ] ]
Image Encoder channels $[128,128]$
learning rate $6*10^{-4}$
Table 2: The key hyperparameters of baselines.
@@
Parameter Settings. Additionally, we provide the detailed experimental hyperparameter settings shown in Table 3 and Table 4, and the description of all the neural networks used in our framework is shown in Appendix C.
Description Value
learning rate $3*10^{-4}$
contrastive batch size $64$
negative pairs number $16$
task embedding size $5$
$\lambda_{r}$ $0.9$
$\lambda_{s}$ $1.0$
encoder hidden size $[256,256,128,128,32]$
decoder hidden size $[64,128]$
Table 3: The hyperparameters of the module for the multiple-encoder and individual-predictor regime.
Description Value
optimizer Adam
$\alpha$ $10^{-4}$
$\beta_{1}$ $0.9$
$\beta_{2}$ $0.999$
$\varepsilon$ $10^{-5}$
seed $[0,10)$
number of process $64$
total frames $10M$
learning rate $10^{-4}$
batch size $1280$
eval interval $1$
eval episodes $500$
image embedding $128$
memory of LSTM $128$
query hidden size $[32,16,8]$
number of head $1$
discounted factor of UCB $0.9$
$m$-step predictor $5$
Table 4: The hyperparameters of the module for the top-K subtask planning tree.
@@
Figure 6: The t-SNE visualizations of subtask embeddings. (a-c): The results are generated by our multiple encoders and one predictor learning regime on 2, 3, and 4 subtasks, respectively. (d-f): The visualization for one shared encoder and one decoder framework.
5.2 Results on Subtask Representation
We present visualization results in Fig. 6 to validate the effectiveness of our subtask representations described in Section 4.2. Specifically, we compare our approach with an ablated version where a shared encoder is used instead of multiple encoders for each subtask. We train both models on 2, 3, and 4 subtasks, respectively, and plot the t-SNE results of subtask embeddings in Fig. 6. The results in Fig. 6(a-c) demonstrate that our approach can clearly differentiate separate clusters of embeddings when using 2, 3, and 4 encoders for each subtask. However, Fig. 6(d-f) shows that using only one shared encoder cannot effectively learn the boundaries and distinguish subtasks. We conjecture that training subtasks with mixed representations may interfere with the stabilization of the representation module.
Moreover, due to the powerful capacity of contrastive learning and model-based dynamic prediction, our model can exhibit strong representational discrimination and contain enough dynamic knowledge, which will be verified empirically as follows.